One of two methods must be used to produce expansion anchors. Method A costs $55,000 initially and will have a $6,000 salvage value after 3 years. The operating cost with this method will be $30,000 per year. Method B will have a first cost of $110,000, an operating cost of $6,000 per year, and a $39,000 salvage value after its 3-year life. The interest rate for both the methods is 10%.

Which method should be used on the basis of a present worth analysis?

The present worth of method A is $  and that of method B is $  .

Method  (Click to select)  A  B  is selected by the company.
Answer #1

NPW of first method = -55000 + 6000 * (P/F, 10%,3) - 30000 * (P/A, 10%,3)

= -55000 + 6000 * 0.751315 - 30000 * 2.486852

= -125097.67 ~ - 125098

NPW of second method = -110000 + 39000 * (P/F, 10%,3) - 6000 * (P/A, 10%,3)

= -110000 + 39000 * 0.751315 - 6000 * 2.486852

= -95619.83 ~ - 95620

As the present cost of second method is less, it should be selected